Gucchi:

Which of the following possibilities will form a triangle? Side = 16 cm, side = 8 cm, side = 7 cm Side = 16 cm, side = 9 cm, side = 7 cm Side = 17 cm, side = 12 cm, side = 7 cm Side = 17 cm, side = 11 cm, side = 5 cm

Gucchi:

I think this is C

Constipated:

@xiogonz

Constipated:

where is your work

Constipated:

how do you know without solving

Constipated:

@XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Gucchi I think this is C \(\color{#0cbb34}{\text{End of Quote}}\) You are Correct

Gucchi:

Alright, Thanks!

Gucchi:

@jimthompson5910 can you give an equation on how it is C?

Sweetie145:

C i think not sure tho

jimthompson5910:

In order for a triangle to be possible, the sum of any two sides must be larger than the third side. In choice A, we have side2+side3 = 8+7 = 15, but this is not larger than side1 = 16. So we rule out choice A. In choice B, we have side2+side3 = 9+7 = 16, but this is also not larger than side1 = 16. We rule out choice B. I'll skip over choice C and come back In choice D, we have side2+side3 = 11+5 = 16 but that's not larger than side1 = 17. So we rule out choice D.

Gucchi:

Wow, This is great! Thanks so much for this.

jimthompson5910:

Now let's focus on choice C only Let's say the three sides are x,y,z x = 17 y = 12 z = 7 note how x+y = 17+12 = 29 which is larger than z = 7 Also, x+z = 17+7 = 24 which is larger than y = 12 Finally, y+z = 12+7 = 19 which is larger than x = 17 So if we pick any two sides of this triangle, add them up, then that sum will be larger than the third side we didn't pick on. This is one example of the triangle inequality theorem and shows why choice C is a possible triangle we can form.

jimthompson5910:

I recommend cutting out slips of paper with the lengths mentioned. Try to physically form triangles based on those lengths. You should find that choices A,B, and D aren't possible triangles.

Gucchi:

True.

Gucchi:

Thank you.

jimthompson5910:

No problem